calculate the molarity of a kcl solution made by adding 15.00ml of distilled water to 48.00ml of 0.75m kcl solution
1
Expert's answer
2020-04-10T13:36:42-0400

С(KCl)= n(KCl)/Vsolution;

n(KCl)= 0.048*0.75=0.036 mol;

Vsolution= 0.048+0.015=0.063 L

C(KCl)=0.036/0.063 = 0.57 M

Answer: the molarity of a kcl solution is 0.57 M